Two trains start from stations 600 km apart and travel toward each other. The trains meet after 6 hours. Find the rate of each train if one travels 10 km per hour faster than the other. can u help me with this problem? thank u.
Answer by Alan3354(30993)

Two trains start from stations 600 km apart and travel toward each other. The trains meet after 6 hours. Find the rate of each train if one travels 10 km per hour faster than the other.
-------------
d = r*t
r = d/t = 600/6 = 100 km/hr
The sum of their speeds is 100 km/hr, the difference is 10 km/hr
x + y = 100
x - y = 10
------------ Add